Molecules in a gas are at 18 degrees celsius. If the rms velocity of the molecules was increased by 4.7%, what temperature would the molecules be raised to?
1
Expert's answer
2020-09-29T09:42:33-0400

"V"rms"= \sqrt\frac{3RT}{M}\times \sqrt T"


"= \alpha T"1/2 "(\alpha = \sqrt\frac{3R}{M})"


"\Delta V"rms"= \frac{4.7 Vrms}{100}"


"\frac{4.7}{100} = 2 \times T\times \frac{4.7}{100}"


"\Delta T = \frac{9.4 T}{100}" -----------------------------1


"\Delta T = \frac{9.4\times 291.5}{100}"


"\Delta T =27.361" ----------------------------2


Final temperature will be

"T"f "= T"2"+ \Delta T"


"= 18+27.368\n\n=45.368 C"